OpenStudy (anonymous):

a 25 gram sample of a substance that's used for drug research has a k-value of 0.1205 Find the substance half life, in days. Round your answer to the nearest tenth

OpenStudy (perl):

$$\Large{ A(t) =A_o e^{kt} \\~\\ A(t) = 25 e^{0.1205 t } } $$

OpenStudy (perl):

A(t) is the amount at t days, and Ao is the initial amount

OpenStudy (perl):

now you want to find the half-life, which is the time it takes to get half the initial amount . so you need to solve $$\Large{ 25/2 = 25 e^{0.1205 t } } $$

OpenStudy (anonymous):

How do I enter an e into the calculator?

OpenStudy (anonymous):

5=25e^0.1205t?

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

-0.693= 0.1205 ?

OpenStudy (perl):

ok k should be negative, since we doing radioactive decay

OpenStudy (perl):

$$\Large{ \frac{25}{2} = 25 e^{-0.1205 t } } $$divide both sides by 25 $$\Large{ \frac{1}{2} = e^{-0.1205 t } } $$now take ln of both sides $$ \Large{ \ln\left(\frac{1}{2}\right) = \ln (e^{-0.1205 t }) } $$use properties of logs, we can bring down exponent $$ \Large{ \ln\left(\frac{1}{2}\right) = -0.1205 t \cdot \ln (e) } $$ ln(e) = 1 $$ \Large{ \ln\left(\frac{1}{2}\right) = -0.1205 t \cdot 1 \\ \frac{\ln\left(\frac{1}{2}\right)}{-0.1205} = t } $$